OpenStudy (anonymous):

help me simplify: sqrt(7) { sqrt(14) + sqrt(3) }

OpenStudy (anonymous):

answer is sqrt(98) + sqart(21)

OpenStudy (anonymous):

you just need to multiply each one of in the bar

OpenStudy (anonymous):

It should look like this \[\sqrt{7} (\sqrt{14} + \sqrt{3}\])

OpenStudy (anonymous):

answer is 7sqrt(2)+sqrt(21)

OpenStudy (anonymous):

thanks